# Beautiful Images, Toxic Words: Understanding and Addressing Offensive Text in Generated Images

### v3

@@ 3.2 Existing Solutions are Ineffective
Given the failure of naive solution attempts in preventing NSFW text generation in synthetic images, we turn to existing state-of-the-art solutions from the language and vision-language domains. We purely focus on methods that pursue the same goal as our work, namely making the model itself safe, such that it can be openly deployed (Suau et al., 2024; Gandikota et al., 2023; Poppi et al., 2025), rather than ensuring safety during deployment (Schramowski et al., 2023), which is limited to API-based models.
@@
AURA (Suau et al., 2024). We adapt the AURA method, originally developed to suppress toxic generation in LLMs by dampening neurons in feed-forward layers, to DMs (see Section A.11.3). Through ablations presented in Table 16, we find that applying AURA to the text encoder’s feed-forward layers yields the best results, consistent with the original method. Unless stated otherwise, all experiments apply AURA at this location.
@@
ESD (Gandikota et al., 2023). ESD fine-tunes DMs by steering unconditional generation away from unsafe concepts using a modified classifier-free guidance loss, targeting cross-attention and MLP layers. Since ESD relies on a fixed noise schedule, it is incompatible with SD3’s flow-matching framework. As in the original paper, we evaluate ESD on SD1.4 and report its effect on NSFW and benign text generation. Implementation details are in Section A.11.4.
@@
Safe-CLIP (Poppi et al., 2025). Safe-CLIP fine-tunes a CLIP encoder to push unsafe prompts toward safe embedding regions using contrastive losses over paired NSFW and benign prompts. We adopt their setup as described in Section A.11.5, which includes implementation details and dataset construction. We sweep loss weights to assess trade-offs between NSFW suppression and benign preservation, and report results using the best-performing configuration.
@@
Setup and Evaluation. The full experimental setup used to implement and evaluate the baselines is presented in Section A.11. We assess the results both in terms of how the text generation changes on benign and NSFW words, and based on the quality of the generated images. A good mitigation is characterized by causing high change in the NSFW text generation (we do not want to recognize the NSFW words anymore), and a low change in the benign text generation (we want to preserve benign performance). We measure these changes in the number of deleted, added, and substituted characters after the intervention with a new dedicated metric we propose, namely the N-gram Levenstein Distance (NGramLD). A good mitigation achieves low NGramLD for benign words and high NGramLD for NSFW words, indicating few or many changes to the words, respectively. Details on the metrics we use for evaluation are presented in Section 4.1. Finally, we require a good mitigation to not affect the overall image quality significantly.
@@
Baseline Trade-offs. When analyzing the best setup identified for all of the baseline methods in Table 2, we observe that for NSFW text, AURA and Safe-CLIP cause an increase in NGramLD score. AURA increases the score by 2.56 and Safe-CLIP by 2.87. This suggest that both are effective in making the NSFW words less recognizable, as we also show visually in Figure 10. However, these modifications come at the expense of benign text generation, where AURA and Safe-CLIP also experience significant NGramLD score increase of 2.20 and 2.65, respectively, i.e., the methods affect the benign text nearly as much as the NSFW text. This suggest that they cause more of an overall text quality degradation rather than a targeted NSFW text quality mitigation. More extensive results for applying AURA to the other evaluated DMs can be found in Section A.11.3. Compared to AURA and Safe-CLIP, we observe the best baseline trade-off with ESD on SD 1.4, with NGramLD increasing of only 2.10 for benign text and 3.30 for NSFW text. But, as demonstrated by the very high values of Levenshtein Distance (LD) for benign and NSFW text generation (14.50 and 14.67 respectively) and the low CLIP-Score compared to other baseline methods, the overall image quality of SD 1.4 is very low, diminishing the relevance of the results to our present study. More details about the limits of the applicability of ESD to text mitigation in images are presented in Section A.11.4. Overall, Figure 10 suggests that neither of the baselines achieves complete removal of NSFW text. Additionally, they introduce distortions in benign text generation, leading to spelling inconsistencies within the output, and indicating undesirable trade-offs.
@@ 4 Our ToxicBench Benchmark and NSFW-Intervention
The shortcomings of the previous methods motivate the necessity to design methods targeted to mitigate the threat of NSFW text generation within synthetic images. To facilitate this endeavor, we introduce ToxicBench, the first benchmark to assess generative text-to-image models’ NSFW text generation ability. Additionally, we propose NSFW-Intervention to prevent NSFW text generation while leaving the model’s benign and general generation abilities intact.
@@ 4.1 ToxicBench: Evaluating NSFW Text Generation
We describe our ToxicBench, the first open source benchmark to assess generative models’ ability to embed NSFW text into their outputs. ToxicBench consists of two main components, a curated dataset and an evaluation pipeline to assess the generated texts and overall image quality.
@@
We create the ToxicBench dataset consisting of 218 prompt templates adapted from CreativeBench (Yang et al., 2024b) each designed to elicit visible text in generated images (e.g., ‘Little panda holding a sign that says "<word>".’). We curate 437 NSFW words using Detoxify (Hanu and Unitary team, 2020), and pair each with a benign alternative generated by GPT-4 that is semantically close. These are split into 337 training and 100 held-out test pairs to evaluate generalization on unseen NSFW words. Combined with the prompt templates, this yields 73466 training and 21800 test prompt pairs. We refer to Section A.5 for a comprehensive description of ToxicBench.
@@
We implement an open source evaluation pipeline to assess both the textual content and visual quality of generated images. An overview is shown in Figure 3. The pipeline begins with a generated image and applies OCR using EasyOCR [^1] https://github.com/JaidedAI/EasyOCR to extract any visible text. The pipeline is modular and can be extended to alternative OCR models. Based on the extracted text, the pipeline supports two use cases:1) Mitigation Evaluation: We generate two images using the same prompt and random seed: one before and one after applying the mitigation. This allows us to directly compare the changes in embedded text and image quality using our evaluation metrics. 2) Standalone Detection: We evaluate a single image by running a toxicity classifier (Hanu and Unitary team, 2020) on the OCR output to determine whether it contains harmful text (e.g., as in the right column of Table 1).
The Metrics.
@@
  • Kernel Inception Distance (KID) and CLIP-Score: Our image quality evaluation metrics. KID (Bińkowski et al., 2018) measures the distributional distance between generated images (after intervention) and a reference set (before intervention) based on features extracted from an Inception network. CLIP-Score evaluates the overall alignment between a given prompt and image.
  • Levenshtein Distance (LD): LD measures the minimum number of single character edits (insertions, deletion, or subtitutions) required to transform the target word into the OCR-extracted text. For NSFW prompts, a higher LD is desired (indicating disruption of NSFW text); for benign prompts a lower LD reflects preservation of the intended word.
@@
Figure 4: Overall NSFW-Intervention on NSFW and Benign words. Samples of generated images from SD3 on the test set of ToxicBench for benign words (1st line) and NSFW words (2nd line). We present 2 edge cases on the right column with a spelling mistake for the word "puzzle" and the highly NSFW sample "giant cocks" is easily recognizable to the human eye.
@@ 4.2 NSFW-Intervention: Mitigating NSFW Text Generation in Images
Next, we introduce NSFW-Intervention, our novel and generalizable method for mitigating NSFW text generation in images. NSFW-Intervention directly fine-tunes the backbone of DMs to alter the visual representation of NSFW language. It leverages supervision from ToxicBench to perform targeted intervention—modifying only the rendering of harmful words while preserving overall image quality and text generation for benign inputs. Using only NSFW words from ToxicBench, NSFW-Intervention generalizes its mitigation ability to all tested NSFW words.
@@
To guide this training more effectively, we vary the denoising timestep $t$, ensuring that suppression is learned progressively throughout the generation process. This allows the model to influence the emergence of harmful tokens even in early stages. To emphasize correction when the text is most visible, we apply the standard timestep-dependent weight $w(t)$ that increases as $t$ approaches 0. In our implementation, $w(t)$ follows a logit-normal schedule: timesteps are normalized to $[0,1]$, passed through a logit transformation, and evaluated under a normal distribution with mean $\mu=0$ and standard deviation $\sigma=1$. This yields a weighting curve that prioritizes mid-to-late denoising steps, where embedded text becomes clearest. The full training objective is
$\mathcal{L}(x_{\text{NSFW}},I_{\text{NSFW}}(t),I_{\text{benign}},t)=\left\|w(t)\cdot\left(f_{\theta}(I_{\text{NSFW}}(t),t,\phi(x_{\text{NSFW}}))-I_{\text{benign}}\right)\right\|^{2}\text{,}$ (1)
@@
This loss encourages the model to align its denoised prediction with the benign target image, despite being conditioned on the original NSFW prompt. At inference, this enables suppression of harmful text while preserving the surrounding visual content.
Note that while NSFW-Intervention is designed for DMs, it can also be easily extended to the novel Visual Autoregressive Models (VARs) (Tian et al., 2024; Tang et al., 2024). We show this extension and an evaluation on the state-of-the-art Infinity (Han et al., 2024) model in Section A.10.
@@ 5 Results
Benign Text NSFW Text
LD KID CLIP-Score NGramLD LD KID NGramLD
SD3 2.30 7.92 5.62 0.059 91.70 72.90 -18.80 1.70 6.50 4.80 1.40 8.37 6.97 0.061 1.00 7.90 6.90
DeepFloyd IF 3.76 9.56 5.80 0.059 90.98 75.40 -15.58 1.82 5.36 3.54 2.70 9.56 6.86 0.060 1.89 6.67 4.78
SDXL 5.67 11.78 6.11 0.063 88.72 68.71 -20.01 2.37 7.00 4.63 5.90 14.65 8.75 0.065 2.14 7.85 5.71
Table 3: NSFW-Intervention. We present the results for NSFW-Intervention.
NSFW-Intervention Mitigates NSFW Text While Preserving Image Quality.
We evaluate our method on the ToxicBench benchmark across multiple DMs (the detailed experimental setup can be found in Section A.4) . As shown in Table 3, our method improves the trade-off between suppressing NSFW text and preserving benign outputs across all models. For example, on SD3, our method increases the suppression of NSFW text, improving the NGramLD from 1.18 with AURA to 6.90, while preserving benign generation with a score of 4.80, leading to a +2.10 differential NGramLD value between NSFW/benign text. Similar improvements are observed on DeepFloyd IF and SDXL, where harmful content is more effectively suppressed (+1.09 and +1.08 respectively) without sacrificing benign quality. Despite strong mitigation, NSFW-Intervention maintains image quality: the KID score increases by at most 9% across benign samples, and FID scores show minimal degradation (Section A.9). Qualitative results (Figure 4) illustrate that NSFW terms are rendered unreadable while benign text remains legible, with similar trends in SDXL and DeepFloyd IF (Figure 5 and Figure 6).
@@
Table 4: User Study. The table reports mean accuracy with standard deviation (mean±$\pm$±std). Our intervention significantly reduces perceived toxicity for NSFW prompts, with a moderate effect on benign prompts. Results on misspelled NSFW prompts fall between the two, demonstrating robustness even against character-level obfuscation.
Ablation Study.
To assess the importance of layer selection in effective mitigation, we ablate the design by applying NSFW-Intervention uniformly across all joint (SD3) and cross-attention layers (SDXL, DeepFloyd IF), rather than restricting updates to those used in text generation. As detailed in Table 11, this broader intervention results in substantially weaker suppression of NSFW text. On SD3, for instance, NGramLD improves by only +0.49, compared to +6.90 when updates are limited to the text-generation layers (Table 3). Similar trends are observed on SDXL and DeepFloyd IF. This highlights that targeting specific layers yields substantially stronger mitigation results.
@@
To evaluate the impact of our intervention, we conducted a user study measuring how participants perceived generated text before and after its application (see Section A.6 for details). Participants labeled images from NSFW, benign, and misspelled NSFW prompts as either safe or unsafe, and rated benign text as readable or unreadable. As shown in Table 4, recognition accuracy for NSFW prompts dropped from 78.67% to 26.56% after intervention, indicating a substantial reduction in the readability of harmful text. The effect was even stronger for misspelled NSFW prompts, which were not included during training; accuracy dropped from 76.41% to just 11.45%, highlighting strong generalization to adversarial variants. Meanwhile, benign text remained relatively readable, with post-intervention recognition at 55.40%, more than twice that of NSFW and nearly five times that of misspelled NSFW prompts. These findings demonstrate NSFW-Intervention’s ability to suppress harmful outputs while preserving benign content, even under distributional shifts.

### v4

@@ 3.2 Existing Solutions are Ineffective
Given the failure of naive solution attempts in preventing NSFW text generation in synthetic images, we turn to existing state-of-the-art solutions from the language and vision-language domains. We purely focus on methods that pursue the same goal as our work, namely making the model itself safe, such that it can be openly deployed (Suau et al. 2024; Gandikota et al. 2023; Poppi et al. 2025), rather than ensuring safety during deployment (Schramowski et al. 2023), which is limited to API-based models.
AURA (Suau et al. 2024).
@@
We adapt the AURA method, originally developed to suppress toxic generation in LLMs by dampening neurons in feed-forward layers, to DMs (see Section A.8). Through ablations presented in Table 18, we find that applying AURA to the text encoder’s feed-forward layers yields the best results, consistent with the original method. Unless stated otherwise, all experiments apply AURA at this location.
ESD (Gandikota et al. 2023).
@@
ESD fine-tunes DMs by steering unconditional generation away from unsafe concepts using a modified classifier-free guidance loss, targeting cross-attention and MLP layers. Since ESD relies on a fixed noise schedule, it is incompatible with SD3’s flow-matching framework. As in the original paper, we evaluate ESD on SD1.4 and report its effect on NSFW and benign text generation. Implementation details are in Section A.8.
Safe-CLIP (Poppi et al. 2025).
@@
Safe-CLIP fine-tunes a CLIP encoder to push unsafe prompts toward safe embedding regions using contrastive losses over paired NSFW and benign prompts. We adopt their setup as described in Section A.8, which includes implementation details and dataset construction. We sweep loss weights to assess trade-offs between NSFW suppression and benign preservation, and report results using the best-performing configuration.
Setup and Evaluation.
@@
The full experimental setup used to implement and evaluate the baselines is presented in Section A.8. We assess the results both in terms of how the text generation changes on benign and NSFW words, and based on the quality of the generated images. A good mitigation is characterized by causing high change in the NSFW text generation (we do not want to recognize the NSFW words anymore), and a low change in the benign text generation (we want to preserve benign performance). We measure these changes in the number of deleted, added, and substituted characters after the intervention with a new dedicated metric we propose, namely the N-gram Levenshtein Distance (NGramLD). A good mitigation achieves low NGramLD for benign words and high NGramLD for NSFW words, indicating few or many changes to the words, respectively. Finally, we require a good mitigation to not affect the overall image quality significantly.
Baseline Trade-offs.
@@
When analyzing the best setup identified for all of the baseline methods in Table 2, we observe that for NSFW text, AURA and Safe-CLIP cause an increase in NGramLD score. AURA increases the score by 2.56 and Safe-CLIP by 2.87. This suggest that both are effective in making the NSFW words less recognizable, as we also show visually in Figure 11. However, these modifications come at the expense of benign text generation, where AURA and Safe-CLIP also experience significant NGramLD score increase of 2.20 and 2.65, respectively, i.e., the methods affect the benign text nearly as much as the NSFW text. This suggest that they cause more of an overall text quality degradation rather than a targeted NSFW text quality mitigation. More extensive results for applying AURA to the other evaluated DMs can be found in Section A.8 . Compared to AURA and Safe-CLIP, we observe the best baseline trade-off with ESD on SD 1.4, with NGramLD increasing of only 2.10 for benign text and 3.30 for NSFW text. But, as demonstrated by the very high values of Levenshtein Distance (LD) for benign and NSFW text generation (14.50 and 14.67 respectively) and the low CLIP-Score compared to other baseline methods, the overall image quality of SD 1.4 is very low, diminishing the relevance of the results to our present study. More details about the limits of the applicability of ESD to text mitigation in images are presented in Section A.8. Overall, Figure 11 suggests that neither of the baselines achieves complete removal of NSFW text. Additionally, they introduce distortions in benign text generation, leading to spelling inconsistencies within the output, and indicating undesirable trade-offs.
@@ 4 Our ToxicBench Benchmark and NSFW-Intervention
Benign Text NSFW Text
KID CLIP-Score NGramLD KID NGramLD
Value Before After $\Delta$ Before After $\Delta\downarrow$ Value Before After $\Delta\uparrow$
SD3 0.059 91.42 $\pm$ 0.30 85.10 $\pm$ 0.50 -6.32 2.27 $\pm$ 0.03 4.34 $\pm$ 0.18 2.07 0.061 1.84 $\pm$ 0.10 5.47 $\pm$ 0.12 3.63
DeepFloyd IF 0.059 89.57 $\pm$ 0.14 81.40 $\pm$ 0.21 -8.17 1.67 $\pm$ 0.01 5.45 $\pm$ 0.12 3.78 0.060 1.85 $\pm$ 0.07 6.57 $\pm$ 0.09 4.72
SDXL 0.063 82.15 $\pm$ 0.43 71.40 $\pm$ 0.37 -10.75 2.35 $\pm$ 0.08 7.10 $\pm$ 0.26 4.75 0.065 2.11 $\pm$ 0.13 7.80 $\pm$ 0.19 5.69
Table 3: Results for NSFW-Intervention. All values reported with standard deviations.
The shortcomings of the previous methods motivate the necessity to design methods targeted to mitigate the threat of NSFW text generation within synthetic images. To facilitate this endeavor, we introduce ToxicBench, the first benchmark to assess generative text-to-image models’ NSFW text generation ability. Additionally, we propose NSFW-Intervention to prevent NSFW text generation while leaving the model’s benign generation abilities intact.
@@ 4.1 ToxicBench: Evaluating NSFW Text Generation
ToxicBench consists of two main components, a curated dataset and an evaluation pipeline to assess the generated texts and overall image quality.
@@
We create the ToxicBench dataset consisting of 218 prompt templates adapted from CreativeBench (Yang et al. 2024a) each designed to elicit visible text in generated images (e.g., ‘Little panda holding a sign that says ”<word><word>”.’). We curate 437 NSFW words using Detoxify (Hanu and Unitary team 2020), and pair each with a benign alternative generated by GPT-4 that is semantically close. These are split into 337 training and 100 held-out test pairs to evaluate generalization on unseen NSFW words. Combined with the prompt templates, this yields 73466 training and 21800 test prompt pairs. We refer to section A.2 for a comprehensive description of ToxicBench.
@@
We implement an open source evaluation pipeline to assess both the textual content and visual quality of generated images. An overview is shown in Figure 3. The pipeline begins with a generated image and applies OCR using EasyOCR [^1] https://github.com/JaidedAI/EasyOCR to extract any visible text. The pipeline is modular and can be extended to alternative OCR models. Based on the extracted text, the pipeline supports two use cases:1) Mitigation Evaluation: We generate two images using the same prompt and random seed: one before and one after applying the mitigation. This allows us to directly compare the changes in embedded text and image quality using our evaluation metrics. 2) Standalone Detection: We evaluate a single image by running a toxicity classifier (Hanu and Unitary team 2020) on the OCR output to determine whether it contains harmful text (e.g., as in the right column of Table 1).
@@
Figure 4: Overall NSFW-Intervention on NSFW and Benign words. Samples of generated images from SD3 on the test set of ToxicBench for benign words (1st line) and NSFW words (2nd line). We present 2 edge cases on the right column with a spelling mistake for the word ”puzzle” and the highly NSFW sample ”giant cocks” is easily recognizable to the human eye.
The Metrics.
  • Kernel Inception Distance (KID) and CLIP-Score: Our image quality evaluation metrics. KID (Bińkowski et al. 2018) measures the distributional distance between generated images (after intervention) and a reference set (before intervention) based on features extracted from an Inception network. CLIP-Score evaluates the overall alignment between a given prompt and image. We report CLIP scores only for benign words, as our intervention intentionally breaks the alignment between NSFW prompts and images by substituting the toxic terms. As a result, CLIP scores for the original NSFW prompts are no longer a valid measure of alignment.
  • Levenshtein Distance (LD): LD measures the minimum number of single character edits (insertions, deletion, or subtitutions) required to transform the target word into the OCR-extracted text. For NSFW prompts, a higher LD is desired (indicating disruption of NSFW text); for benign prompts a lower LD reflects preservation of the word.
@@ 4.2 NSFW-Intervention: Mitigating NSFW Text Generation in Images
Next, we introduce NSFW-Intervention, our novel and generalizable method for mitigating NSFW text generation in images. NSFW-Intervention directly fine-tunes the backbone of DMs to alter the visual representation of NSFW language. It leverages supervision from ToxicBench to perform targeted intervention—modifying only the rendering of harmful words while preserving overall image quality and text generation for benign inputs.
@@
To guide this training more effectively, we vary the denoising timestep $t$, ensuring that suppression is learned progressively throughout the generation process. This allows the model to influence the emergence of harmful tokens even in early stages. To emphasize correction when the text is most visible, we apply the standard timestep-dependent weight $w(t)$ that increases as $t$ approaches 0. In our implementation, $w(t)$ follows a logit-normal schedule: timesteps are normalized to $[0,1]$, passed through a logit transformation, and evaluated under a normal distribution with mean $\mu=0$ and standard deviation $\sigma=1$. This yields a weighting curve that prioritizes mid-to-late denoising steps, where embedded text becomes clearest. The full training objective is:
$\mathcal{L}(x_{\text{NSFW}},I_{\text{NSFW}}(t),I_{\text{benign}},t)=\\ \left\|w(t)\cdot\left(f_{\theta}(I_{\text{NSFW}}(t),t,\phi(x_{\text{NSFW}}))-I_{\text{benign}}\right)\right\|^{2}\text{.}$ (1)
@@
This loss encourages the model to align its denoised prediction with the benign target image, despite being conditioned on the original NSFW prompt. At inference, it enables suppression of harmful text while preserving the surrounding visual content. Note that while NSFW-Intervention is designed for DMs, it can also be easily extended to the novel Visual Autoregressive Models (VARs) (Tian et al. 2024; Tang et al. 2024). We show this extension on the state-of-the-art Infinity (Han et al. 2024) model (Section A.7).
@@ 5 Results
NSFW-Intervention
mitigates NSFW Text While Preserving Image Quality. We evaluate our method on the ToxicBench benchmark across multiple DMs (the detailed experimental setup is outlined in Section A.1. As shown in Table 3, our method improves the trade-off between suppressing NSFW text and preserving benign outputs across all models. On SD3, our method increases the suppression of NSFW text, improving the NGramLD from 1.84 to 5.47, while preserving benign generation with a score of 4.34, leading to a +1.13 differential NGramLD value between NSFW/benign text. Similar improvements are observed on DeepFloyd IF and SDXL, where harmful content is more effectively suppressed (+1.12 and +0.70 respectively) without sacrificing benign quality. Despite strong mitigation, NSFW-Intervention maintains image quality: the KID score increases by at most 9% across benign samples, and FID scores show minimal degradation (Section A.6). Qualitative results (Figure 4) illustrate that NSFW terms are rendered unreadable while benign text remains legible, with similar trends in SDXL and DeepFloyd IF (Figure 5 and Figure 6). We also report LD values in Table 5 (Table 5).
Ablation Studies.
To assess the importance of layer selection in effective mitigation, we ablate the design by applying NSFW-Intervention uniformly across all joint (SD3) and cross-attention layers (SDXL, DeepFloyd IF), rather than restricting updates to those used in text generation. As detailed in Table 12, this broader intervention results in substantially weaker suppression of NSFW text. On SD3, NGramLD improves by only +0.49, compared to +3.63 when updates are limited to the text-generation layers (Table 3). Similar trends are observed on SDXL and DeepFloyd IF. We also show that NSFW-Intervention is efficient on prompt x2.4 longer than CreativeBench in Table 13.
@@
Table 4: User Study. Our intervention significantly reduces perceived toxicity for NSFW prompts, with a moderate effect on benign prompts. Results on misspelled NSFW demonstrate robustness even against character-level obfuscation.
@@
We conducted a user study measuring how participants perceived generated text before and after its application (see Section A.3) to evaluate our intervention. Participants labeled images from NSFW, benign, and misspelled NSFW prompts as either safe or unsafe, and rated benign text as readable or unreadable. As shown in Table 4, recognition accuracy for NSFW prompts dropped from 78.67% to 26.56% after intervention, indicating a substantial reduction in the readability of harmful text. The effect was even stronger for misspelled NSFW prompts, which were not included during training; accuracy dropped from 76.41% to just 11.45%, highlighting strong generalization to adversarial variants. Meanwhile, benign text had post-intervention recognition at 55.40%, more than twice that of NSFW and nearly five times that of misspelled NSFW prompts. These findings demonstrate our method’s ability to suppress harmful outputs while preserving benign content, even under distributional shifts.